OpenStudy (anonymous):

solve for variables in the equation below 12^(y+2)=6 e^x=11 Round your answers to the nearest hundredth. Do not round any intermediate computations.

OpenStudy (anonymous):

@pkamania

OpenStudy (mathmate):

1. Take log to (base 12) on both sides to get log(base 12)12^(y+2)=log(base 12)6=log(6)/log(12) y+2 = log(6)/log(12) Solve for y. 2. Same as 1, but take log to base e (natural log) log(base e)e^x = log(base e) 11 x = log(base e) 11